# MBA-SLAM: Motion Blur Aware Dense Visual SLAM with Radiance Fields Representation

### v1

@@ Abstract
Emerging 3D scene representations, such as Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS), have demonstrated their effectiveness in Simultaneous Localization and Mapping (SLAM) for photo-realistic rendering, particularly when using high-quality video sequences as input. However, existing methods struggle with motion-blurred frames, which are common in real-world scenarios like low-light or long-exposure conditions. This often results in a significant reduction in both camera localization accuracy and map reconstruction quality. To address this challenge, we propose a dense visual SLAM pipeline (i.e. MBA-SLAM) to handle severe motion-blurred inputs. Our approach integrates an efficient motion blur-aware tracker with either neural radiance fields or Gaussian Splatting based mapper. By accurately modeling the physical image formation process of motion-blurred images, our method simultaneously learns 3D scene representation and estimates the cameras’ local trajectory during exposure time, enabling proactive compensation for motion blur caused by camera movement. In our experiments, we demonstrate that MBA-SLAM surpasses previous state-of-the-art methods in both camera localization and map reconstruction, showcasing superior performance across a range of datasets, including synthetic and real datasets featuring sharp images as well as those affected by motion blur, highlighting the versatility and robustness of our approach. Code: https://github.com/WU-CVGL/MBA-SLAM.
@@
Simultaneous Localization and Mapping, Neural Radiance Fields, 3D Gaussian Splatting
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ 2 Related Work
Several scene deblurring methods based on NeRF [7] and 3DGS [8] have emerged, such as Deblur-NeRF [28], DP-NeRF [29], Deblur-GS [30] and BAGS [31]. These methods aim to reconstruct sharp scene representations from sets of motion-blurred images by estimating the blur kernel while fixing the inaccurate camera poses recovered from blurred images during training NeRF and 3DGS. BAD-NeRF [22] and its extension, ExBlurF [32], BeNeRF [33] and BAD-Gaussians [23], employ joint learning of camera motion trajectories within exposure time and radiance fields, adhering to the physical blur process. Despite these advancements, the accurate initialization of camera poses from COLMAP [34] remains a prerequisite, and challenges pertaining to low convergence efficiency persist, hindering their seamless integration into SLAM systems.
@@
Figure 1: The pipeline of MBA-SLAM. Our framework consists of blur aware tracking process and bundle adjustment deblurring mapping process. Tracking: Given the current blurry frame, the mapper first renders a virtual sharp image of the lastest blurry keyframe from the 3D scene. Our motion blur-aware tracker directly estimates the camera motion trajectory during the exposure time, represented by the camera positions at the start and end of the exposure (i.e.$\mathbf{T}_{\mathrm{start}}$ and $\mathbf{T}_{\mathrm{end}}$). Intermediate camera poses can be interpolated in $\mathbf{SE}(3)$ space. Mapping: Our mapper generates virtual sharp images along the camera trajectory, following the standard rendering procedures of Radiance Fields or Gaussian Splatting. The blurry image can then be synthesized by averaging these virtual images, adhering to the physical image formation model of motion-blurred images. Finally, the scene representation and camera trajectory are jointly optimized by minimizing the loss between the synthesized images and the input data.
@@ 3.2.2 Camera Motion Trajectory modeling
The Jacobians, i.e.$\frac{\partial\mathbf{t}_{t}}{\partial\mathbf{t}_{\mathrm{start}}}\in\mathbb{R}^{3\times 3}$ and $\frac{\partial\mathbf{t}_{t}}{\partial\mathbf{t}_{\mathrm{end}}}\in\mathbb{R}^{3\times 3}$ can thus be derived as
@@ 3.2.3 Direct Image Alignment with Blurry Images
Figure 2: Pixel point transfer strategies. Note that we assume the pixel center lies at the grid intersection, e.g.the green grid is considered as a 3$\times$3 patch.
Each sampled pixel in $\mathbf{I}_{\mathrm{ref}}$ with a known depth is transfered to the current (blurry) image $\mathbf{B}_{\mathrm{cur}}$ using (4). We then identify the nearest integer position pixel in the current blurry image for each projected point. Assuming the 3D point lies on a fronto-parallel plane (with respect to $\mathbf{I}_{\mathrm{ref}}$), we leverage this plane to transfer the selected pixel back into the reference view. Further details are provided in Fig. 2. To synthesize the re-blurred pixel from the reference view (facilitating comparison with the real captured pixel intensity), we interpolate between $\mathbf{T}_{\mathrm{start}}$ and $\mathbf{T}_{\mathrm{end}}$. For each virtual view $\mathbf{T}_{t}$ sampled uniformly within $[0,\tau]$, the pixel coordinate (i.e.the red pixel in Fig. 2) is transferred back into the reference image and image intensity values are obtained by bi-linear interpolation. The re-blurred pixel intensity is then computed by averaging over the intensity values (as in (2)):
@@ 3.2.4 More Details on the Transfer
We will illustrate the relationship in Fig. 3 to mitigate potential confusion. We denote the translation vector $\mathbf{p}_{i}^{ref}$ of $\mathbf{T}_{i}^{ref}$ with $[p_{x},p_{y},p_{z}]^{T}$ and represent the rotation $\mathbf{R}_{i}^{ref}$ with unit quaternion, i.e.$\bar{\mathbf{q}}$ = $[q_{x}$, $q_{y}$, $q_{z}$, $q_{w}]^{T}$. We denote $d$ as the depth of the frontal-parallel plane with respect to the reference key-frame. We can then compute the distance $d^{\prime}$ between the camera center of the $i^{th}$ virtual camera to the frontal-parallel plane as
@@
where $x^{2}+y^{2}+z^{2}=1$. We can then compute cosine function of the angle (i.e.$\theta$) between the unitary ray and the plane normal of the frontal parallel plane as
@@
Jacobian derivations: The pose of the $i^{th}$ virtual camera, i.e.$\mathbf{T}_{i}^{ref}$, relates to $\mathbf{T}_{0}^{ref}$ and $\mathbf{T}_{\tau}^{ref}$ via Eq. (5). To estimate both $\mathbf{T}_{0}^{ref}$ and $\mathbf{T}_{\tau}^{ref}$, we need to have the Jacobian of $\mathbf{x}_{\frac{i\tau}{n-1}}$ with respect to $\mathbf{T}_{i}^{ref}$. Since the relationship between $\mathbf{x}_{\frac{i\tau}{n-1}}$ and $\mathbf{T}_{i}^{ref}$ is complex, as derived above, we use the Mathematica Symbolic Toolbox [^3] https://www.wolfram.com/mathematica/ for the ease of Jacobian derivations. The details are as follows.
@@ 4.1 Experimental Setup
Metrics. We utilize the widely adopted absolute trajectory error (i.e.ATE RMSE) [21] to assess tracking performance. In evaluating reconstruction and geometric accuracy, we employ the culled mesh [52] to compute the 2D Depth L1 metric (cm), as well as 3D metrics such as Precision (%), Recall (%) and F1-score ($<$ 1cm %). To fairly compare these baselines, we run all the methods 5 times and report the average results. We further evaluate the rendering quality using classical signal-to-noise ratio (PSNR), SSIM and LPIPS [53] metrics.
Implementation Details.  Tracker: To enhance efficiency, we sub-sample high-gradient pixels to derive sparse keypoints, ensuring uniform distribution across the image. Additionally, we define $9\times 9$ local patches around each sampled sparse keypoint to facilitate better convergence. The energy function undergoes optimization in a coarse-to-fine manner, supplemented by the application of a robust huber loss function for improved robustness. The CUDA-accelerated backpropagation process implemented in our tracker is well-suited for real-time applications.  Mapper: We implement the mapper using either Radiance Fields or 3DGS. If mapper is Radiance Fields: We interpolate 7 virtual images between $\mathbf{T}_{\mathrm{start}}$ and $\mathbf{T}_{\mathrm{end}}$ ($n$ in Eq. 2) following [22]. Other details are same as [12]. If mapper is Gaussian Splatting: We implement the GS version of MBA-SLAM using PyTorch within the gsplat framework [54]. The number of virtual images is set to 13. The coefficients and regularization $r$ for $\mathcal{L}_{reg}$ is $11e$ 1 and 1.0. Other configurations are as default in gsplat. As for Gaussians management, we add gaussians for new frame at those areas that alpha value is smaller than 0.5 and we adopt the default splitting, cloning, and pruning strategy in gsplat. For Gaussian color initialization, we directly utilize input blur color and depth information on these pixels to spawn newly added Gaussian primitives. Additionally, we have created interfaces in the tracker to facilitate PyTorch data transmission from the mapper to our CUDA-implemented tracker.
Input Refer to caption CoSLAM [11] Refer to caption
Point SLAM [26] Refer to caption ESLAM [12] Refer to caption
Spla TAM [15] Refer to caption Ours-NeRF Refer to caption
Groundtruth Refer to caption Ours-GS Refer to caption
@@
NICESL AM [10] PSNR $\uparrow$ - Precision $\uparrow$
@@
VoxFus ion [51] PSNR $\uparrow$ 17.97 17.07 13.52 16.19 Precision $\uparrow$ 38.01
@@
CoSLA [11] PSNR $\uparrow$ 22.97 21.98 21.45 22.13 Precision $\uparrow$ 55.71
@@
ESL AM [12] PSNR $\uparrow$ 21.07 23.87 - Precision $\uparrow$ 44.61
@@
Point SLAM [26] PSNR $\uparrow$ 15.90 18.29 - Precision $\uparrow$ 1.45
@@
Ours NeRF PSNR $\uparrow$ 24.77 26.35 24.13 25.08 Precision $\uparrow$ 82.49
@@
Spla TAM [15] PSNR $\uparrow$ 17.69 - Precision $\uparrow$ 15.57
@@
Ours-GS PSNR $\uparrow$ 28.45 30.16 27.85 28.82 Precision $\uparrow$ 83.71
@@
Figure 6: Qualitative mesh visulization of different methods with ArchViz-1 datasets. The result reveals that implicit Radiance Fields (e.g.CoSLAM, ESLAM) deliver better reconstruction mesh performance than explicit point based methods (i.e.Point-SLAM, SplaTAM). MBA-SLAM always achieves best performance, no matter Ours-NeRF or Our-GS.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input CoSLAM [11] ESLAM [12] PointSLAM [26] SplaTAM [15] Ours-GS
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input CoSLAM [11] ESLAM [12] PointSLAM [26] SplaTAM [15] Ours-GS
Figure 8: Qualitative rendering results of different methods with our real captured Realsense datasets. The experimental results demonstrate that our method achieves superior performance over prior methods on the real captured dataset as well.
4.2 Evaluation on Blur Datasets
@@
TABLE IV: Tracking comparison (ATE RMSE [cm]) of the proposed method vs. the SOTA methods on the #Replica dataset. The best and second-best results are colored. MBA-SLAM achieves best tracking performance in Radiance Fields and Gaussian Splatting approaches, respectively.
@@
TABLE V: Rendering Performance on Replica. We outperform existing dense SLAMs on the commonly reported rendering metrics.
@@
NICESL AM [10] PSNR $\uparrow$ 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 24.42
@@
VoxFus ion [51] PSNR $\uparrow$ 22.39 22.36 23.92 27.79 29.83 20.33 23.47 25.21 24.41
@@
CoSL AM [11] PSNR $\uparrow$ 27.27 28.45 29.06 34.14 34.87 28.43 28.76 30.91 30.24
@@
Ours NeRF PSNR $\uparrow$ 28.20 29.84 29.92 35.27 35.03 27.03 29.49 32.01 30.85
@@
GSSL AM [15] PSNR $\uparrow$ 31.56 32.86 32.59 38.70 41.17 32.36 32.03 32.92 34.27
@@
SplaT AM [15] PSNR $\uparrow$ 32.34 33.66 35.34 38.39 39.33 32.18 30.27 32.37 34.24
@@
Ours GS PSNR $\uparrow$ 35.68 34.75 38.63 40.86 42.26 33.26 33.69 36.69 36.98
@@
NICESL AM [10] Depth L1 $\downarrow$ 1.92 1.37 1.83 1.47 1.50 2.68 2.84 1.95 1.95
@@
VoxFus ion [51] Depth L1 $\downarrow$ 0.55 0.63 1.00 2.96 1.30 1.25 1.37 0.76 1.23
@@
CoSLA [11] Depth L1 $\downarrow$ 0.87 0.91 2.26 1.17 1.25 2.13 2.30 1.51 1.55
@@
Ours NeRF Depth L1 $\downarrow$ 0.52 0.37 0.71 0.56 0.92 1.68 1.12 0.51 0.80
@@
Spla TAM [15] Depth L1 $\downarrow$ 0.49 0.37 0.59 0.43 0.63 0.97 1.15 0.71 0.67
@@
Ours GS Depth L1 $\downarrow$ 0.47 0.39 0.56 0.35 0.63 1.46 1.51 0.55 0.74
@@
Figure 9: Qualitative mesh visulization results of different methods with Replica datasets. It demonstrate MBA-SLAM surpass other state-og-the-art dense visual SLAMs even on standard sharp datasets.

### v2

# MBA-SLAM: Motion Blur Aware Gaussian Splatting SLAM

@@ Abstract
Emerging 3D scene representations, such as Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS), have demonstrated their effectiveness in Simultaneous Localization and Mapping (SLAM) for photo-realistic rendering, particularly when using high-quality video sequences as input. However, existing methods struggle with motion-blurred frames, which are common in real-world scenarios like low-light or long-exposure conditions. This often results in a significant reduction in both camera localization accuracy and map reconstruction quality. To address this challenge, we propose a dense visual deblur SLAM pipeline (i.e. MBA-SLAM) to handle severe motion-blurred inputs and enhance image deblurring. Our approach integrates an efficient motion blur-aware tracker with either neural radiance fields or Gaussian Splatting based mapper. By accurately modeling the physical image formation process of motion-blurred images, our method simultaneously learns 3D scene representation and estimates the cameras’ local trajectory during exposure time, enabling proactive compensation for motion blur caused by camera movement. In our experiments, we demonstrate that MBA-SLAM surpasses previous state-of-the-art methods in both camera localization and map reconstruction, showcasing superior performance across a range of datasets, including synthetic and real datasets featuring sharp images as well as those affected by motion blur, highlighting the versatility and robustness of our approach. Code is available at https://github.com/WU-CVGL/MBA-SLAM.
@@
SLAM, Deblur, Neural Radiance Fields, 3D Gaussian Splatting
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ 2 Related Work
Several scene deblurring methods based on NeRF [7] and 3DGS [8] have emerged, such as Deblur-NeRF [28], DP-NeRF [29], Deblur-GS [30] and BAGS [31]. These methods aim to reconstruct sharp scene representations from sets of motion-blurred images by estimating the blur kernel while fixing the inaccurate camera poses recovered from blurred images during training NeRF and 3DGS. BAD-NeRF [22] and its extension, ExBlurF [32] and BAD-Gaussians [23], employ joint learning of camera motion trajectories within exposure time and radiance fields, adhering to the physical blur process. Despite these advancements, the accurate initialization of camera poses from COLMAP [33] remains a prerequisite, and challenges pertaining to low convergence efficiency persist, hindering their seamless integration into SLAM systems.
@@
[…]
@@ 3.2.2 Camera Motion Trajectory modeling
[…]
@@ 3.2.3 Direct Image Alignment with Blurry Images
Figure 2: Pixel point transfer strategies. Note that we assume the pixel center lies at the grid intersection, e.g.the green grid is considered as a $\times3$ patch.
Each sampled pixel in $\mathbf{I}_{\mathrm{ref}}$ with a known depth is transfered to the current (blurry) image $\mathbf{B}_{\mathrm{cur}}$ using (4). We then identify the nearest integer position pixel in the current blurry image for each projected point. Assuming the 3D point lies on a fronto-parallel plane (with respect to $\mathbf{I}_{\mathrm{ref}}$), we leverage this plane to transfer the selected pixel back into the reference view. Further details are provided in Fig. 2. To synthesize the re-blurred pixel from the reference view (facilitating comparison with the real captured pixel intensity), we interpolate between $\mathbf{T}_{\mathrm{start}}$ and $\mathbf{T}_{\mathrm{end}}$. For each virtual view $\mathbf{T}_{t}$ sampled uniformly within $0,\tau]$, the pixel coordinate (i.e.the red pixel in Fig. 2) is transferred back into the reference image and image intensity values are obtained by bi-linear interpolation. The re-blurred pixel intensity is then computed by averaging over the intensity values (as in (2)):
@@ 3.2.4 More Details on the Transfer
[…]
@@
[…]
@@
[…]
@@ 4.1 Experimental Setup
Metrics. We utilize the widely adopted absolute trajectory error (i.e.ATE RMSE) [21] to assess tracking performance. In evaluating reconstruction and geometric accuracy, we employ the culled mesh [52] to compute the 2D Depth L1 metric (cm), as well as 3D metrics such as Precision (%), Recall (%) and F1-score (<< 1cm %). To fairly compare these baselines, we run all the methods 5 times and report the average results. We further evaluate the rendering quality using classical signal-to-noise ratio (PSNR), SSIM and LPIPS [53] metrics.
Implementation Details.  Tracker: To enhance efficiency, we sub-sample high-gradient pixels to derive sparse keypoints, ensuring uniform distribution across the image. Additionally, we define $9\times 9$ local patches around each sampled sparse keypoint to facilitate better convergence. The energy function undergoes optimization in a coarse-to-fine manner, supplemented by the application of a robust huber loss function for improved robustness. The CUDA-accelerated backpropagation process implemented in our tracker is well-suited for real-time applications.  Mapper: We implement the mapper using either Radiance Fields or 3DGS. If mapper is Radiance Fields: We interpolate 7 virtual images between $\mathbf{T}_{\mathrm{start}}$ and $\mathbf{T}_{\mathrm{end}}$ ($n$ in Eq. 2) following [22]. Other details are same as [12]. If mapper is Gaussian Splatting: We implement the GS version of MBA-SLAM using PyTorch within the gsplat framework [54]. The number of virtual images is set to 13. The coefficients and regularization $r$ for $\mathcal{L}_{reg}$ is $1e1$ and 1.0. Other configurations are as default in gsplat. As for Gaussians management, we add gaussians for new frame at those areas that alpha value is smaller than 0.5 and we adopt the default splitting, cloning, and pruning strategy in gsplat. For Gaussian color initialization, we directly utilize input blur color and depth information on these pixels to spawn newly added Gaussian primitives. Additionally, we have created interfaces in the tracker to facilitate PyTorch data transmission from the mapper to our CUDA-implemented tracker.
Input Refer to caption Groundtruth Refer to caption
CoSLAM[11] Refer to caption Spla TAM [15] Refer to caption
ESLAM [12] Refer to caption MonoGS[16] Refer to caption
Point SLAM [26] Refer to caption Photo SLAM [17] Refer to caption
Ours-NeRF Refer to caption Ours-GS Refer to caption
@@
NICESL
AM [10]
PSNR $\uparrow$ - Precision $\uparrow$
@@
VoxFus
ion [50]
PSNR $\uparrow$ 17.97 17.07 13.52 16.19 Precision $\uparrow$ 38.01
@@
CoSLA
[11]
PSNR $\uparrow$ 22.97 21.98 21.45 22.13 Precision $\uparrow$ 55.71
@@
ESL
AM [12]
PSNR $\uparrow$ 21.07 23.87 - Precision $\uparrow$ 44.61
@@
Point
SLAM [26]
PSNR $\uparrow$ 15.90 18.29 - Precision $\uparrow$ 1.45
@@
Ours
NeRF
PSNR $\uparrow$ 24.77 26.35 24.13 25.08 Precision $\uparrow$ 82.49
@@
Spla
TAM [15]
PSNR $\uparrow$ 17.69 - Precision $\uparrow$ 15.57
@@
RTG
SLAM [51]
PSNR $\uparrow$ 20.08 17.99 - Precision $\uparrow$
SSIM $\uparrow$ 0.558 0.595 - Recall $\uparrow$
LPIPS $\downarrow$ 0.551 0.538 - F1 $\uparrow$
Photo
SLAM [17]
PSNR $\uparrow$ 23.45 27.49 23.70 24.88 Precision $\uparrow$ 52.69
SSIM $\uparrow$ 0.823 0.897 0.846 0.855 Recall $\uparrow$ 63.88
LPIPS $\downarrow$ 0.387 0.363 0.396 0.382 F1 $\uparrow$ 53.76
Mono
GS [16]
PSNR $\uparrow$ 23.18 27.14 20.56 23.63 Precision $\uparrow$ 44.43
SSIM $\uparrow$ 0.852 0.884 0.696 0.811 Recall $\uparrow$ 57.61
LPIPS $\downarrow$ 0.337 0.329 0.427 0.364 F1 $\uparrow$ 50.17
Ours-GS
PSNR $\uparrow$ 28.45 30.16 27.85 28.82 Precision $\uparrow$ 83.71
@@
Photo SLAM[17] Refer to caption
MonoGS[16] Refer to caption
@@
Figure 6: Qualitative mesh visualization of different methods with ArchViz-1 datasets. The result reveals that implicit Radiance Fields (e.g.CoSLAM, ESLAM) deliver better reconstruction mesh performance than explicit point based methods (i.e.Point-SLAM, SplaTAM). MBA-SLAM always achieves best performance, no matter Ours-NeRF or Our-GS. RTG-SLAM fails to reconstruct mesh.
@@
4.2 Evaluation on Blur Datasets
Input Refer to caption
CoSLAM[11] Refer to caption
ESLAM[12] Refer to caption
Point SLAM[26] Refer to caption
Spla TAM[15] Refer to caption
RTG SLAM[51] Refer to caption
Photo SLAM[17] Refer to caption
MonoGS[16] Refer to caption
Ours-GS Refer to caption
@@ 4.2 Evaluation on Blur Datasets
Input Refer to caption
CoSLAM[11] Refer to caption
ESLAM[12] Refer to caption
Point SLAM[26] Refer to caption
Spla TAM[15] Refer to caption
RTG SLAM[51] Refer to caption
Photo SLAM[17] Refer to caption
MonoGS[16] Refer to caption
Ours-GS Refer to caption
Figure 8: Qualitative rendering results of different methods with our real captured Realsense datasets. The experimental results demonstrate that our method achieves superior performance over prior methods on the real captured dataset as well. Note that the first column shows an incorrect view rendered by Photo-SLAM, attributed to a failure in tracking.
@@
TABLE IV: Tracking comparison (ATE RMSE [cm]) of the proposed method vs. the SOTA methods on the #Replica dataset. The best and second-best results are colored.
@@
TABLE V: Rendering Performance on Replica. While Point-SLAM outperforms other NeRF-based SLAM methods, it suffers from significantly longer mapping optimization times — over 30 times slower than CoSLAM, ESLAM, and our NeRF-based system, as shown in Table VIII.
@@
NICESL
AM [10]
PSNR $\uparrow$ 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 24.42
@@
VoxFus
ion [50]
PSNR $\uparrow$ 22.39 22.36 23.92 27.79 29.83 20.33 23.47 25.21 24.41
@@
CoSL
AM [11]
PSNR $\uparrow$ 27.27 28.45 29.06 34.14 34.87 28.43 28.76 30.91 30.24
@@
Point
SLAM [26]
PSNR $\uparrow$ 32.40 34.08 35.50 38.26 39.16 33.99 33.48 33.49 35.05
SSIM $\uparrow$ 0.974 0.977 0.982 0.983 0.986 0.960 0.960 0.979 0.975 1.33
LPIPS $\downarrow$ 0.113 0.116 0.111 0.100 0.118 0.156 0.132 0.142 0.124
Ours
NeRF
PSNR $\uparrow$ 28.20 29.84 29.92 35.27 35.03 27.03 29.49 32.01 30.85
@@
GSSL
AM [14]
PSNR $\uparrow$ 31.56 32.86 32.59 38.70 41.17 32.36 32.03 32.92 34.27
@@
SplaT
AM [15]
PSNR $\uparrow$ 32.34 33.66 35.34 38.39 39.33 32.18 30.27 32.37 34.24
@@
RTG
SLAM [51]
PSNR $\uparrow$ 25.53 29.91 31.19 36.65 37.19 29.88 29.96 33.11 31.68
SSIM $\uparrow$ 0.900 0.919 0.937 0.984 0.985 0.958 0.959 0.952 0.949 456.89
LPIPS $\downarrow$ 0.269 0.212 0.193 0.096 0.108 0.187 0.172 0.158 0.174
Photo
SLAM [17]
PSNR $\uparrow$ 31.68 35.50 36.78 38.97 39.78 33.94 34.26 36.77 35.96
SSIM $\uparrow$ 0.948 0.970 0.987 0.984 0.988 0.975 0.973 0.969 0.974 1084.00
LPIPS $\downarrow$ 0.100 0.080 0.061 0.076 0.079 0.126 0.095 0.078 0.087
Mono
GS [16]
PSNR $\uparrow$ 34.55 36.72 37.28 40.06 41.57 35.13 35.71 36.03 37.13
SSIM $\uparrow$ 0.969 0.982 0.989 0.991 0.989 0.990 0.987 0.982 0.985 769.00
LPIPS $\downarrow$ 0.071 0.069 0.071 0.051 0.073 0.073 0.079 0.064 0.069
Ours
GS
PSNR $\uparrow$ 35.68 34.75 38.63 40.86 42.26 33.26 33.69 36.69 36.98
@@
NICESL
AM [10]
Depth L1 $\downarrow$ 1.92 1.37 1.83 1.47 1.50 2.68 2.84 1.95 1.95
@@
VoxFus
ion [50]
Depth L1 $\downarrow$ 0.55 0.63 1.00 2.96 1.30 1.25 1.37 0.76 1.23
@@
CoSLA
[11]
Depth L1 $\downarrow$ 0.87 0.91 2.26 1.17 1.25 2.13 2.30 1.51 1.55
@@
PointSL
AM [26]
Depth L1 $\downarrow$ 0.53 0.22 0.46 0.30 0.57 0.49 0.51 0.46 0.44
Precision $\uparrow$ 91.95 99.04 97.89 99.00 99.37 98.05 96.61 93.98 96.99
Recall $\uparrow$ 82.48 86.43 84.64 89.06 84.99 81.44 81.17 78.51 83.59
F1 $\uparrow$ 86.90 92.31 90.78 93.77 91.62 88.98 88.22 85.55 89.77
Ours
NeRF
Depth L1 $\downarrow$ 0.52 0.37 0.71 0.56 0.92 1.68 1.12 0.51 0.80
@@
Spla
TAM [15]
Depth L1 $\downarrow$ 0.49 0.37 0.59 0.43 0.63 0.97 1.15 0.71 0.67
@@
RTGS
LAM [51]
Depth L1 $\downarrow$ 0.79 0.85 1.49 0.98 1.12 1.97 1.78 0.92 1.24
Precision $\uparrow$ 79.75 83.49 78.79 81.66 82.40 76.44 70.12 77.43 78.76
Recall $\uparrow$ 79.46 75.38 71.52 75.20 71.93 68.34 63.72 69.86 71.93
F1 $\uparrow$ 80.05 79.21 75.25 78.29 75.58 72.66 66.68 72.51 75.03
PhotoSL
AM [17]
Depth L1 $\downarrow$ 0.87 0.41 0.98 0.45 0.72 1.21 1.34 0.64 0.83
Precision $\uparrow$ 85.89 94.20 91.59 93.84 96.04 92.96 88.07 91.63 91.78
Recall $\uparrow$ 81.74 83.71 80.12 87.09 83.24 77.82 75.32 75.86 80.61
F1 $\uparrow$ 82.31 89.23 85.27 91.98 90.16 84.72 82.04 82.45 86.02
Mono
GS [16]
Depth L1 $\downarrow$ 0.69 0.39 0.61 0.55 0.67 1.39 0.96 0.68 0.74
Precision $\uparrow$ 88.24 94.93 95.31 93.14 97.75 94.04 90.17 86.85 92.55
Recall $\uparrow$ 81.56 83.31 82.93 84.35 83.04 78.68 76.09 73.82 80.47
F1 $\uparrow$ 85.40 88.74 88.38 88.53 89.93 86.12 82.86 79.81 86.22
Ours
GS
Depth L1 $\downarrow$ 0.47 0.39 0.56 0.35 0.63 1.46 1.51 0.55 0.74
@@
RTG SLAM[51] Refer to caption
Photo SLAM[17] Refer to caption
MonoGS[16] Refer to caption
@@
Figure 9: Qualitative mesh visualization results of different methods with Replica datasets. It demonstrates MBA-SLAM surpasses other state-of-the-art dense visual SLAMs even on standard sharp datasets. Note that RTG-SLAM produces incomplete meshes due to the presence of numerous holes in the rendered depth maps.